### v1

# Hi-Core: Hierarchical Knowledge Transfer for Continual Reinforcement Learning

Abstract
Continual reinforcement learning (CRL) empowers RL agents with the ability to learn from a sequence of tasks, preserving previous knowledge and leveraging it to facilitate future learning. However, existing methods often focus on transferring low-level knowledge across similar tasks, which neglects the hierarchical structure of human cognitive control, resulting in insufficient knowledge transfer across diverse tasks. To enhance high-level knowledge transfer, we propose a novel framework named Hi-Core (Hierarchical knowledge transfer for Continual reinforcement learning), which is structured in two layers: 1) the high-level policy formulation which utilizes the powerful reasoning ability of the Large Language Model (LLM) to set goals and 2) the low-level policy learning through RL which is oriented by high-level goals. Moreover, the knowledge base (policy library) is constructed to store policies that can be retrieved for hierarchical knowledge transfer. Experiments conducted in MiniGrid have demonstrated the effectiveness of Hi-Core in handling diverse CRL tasks, outperforming popular baselines.
1 Introduction
Reinforcement learning (RL) is a powerful paradigm in artificial intelligence (AI) that enables agents to learn optimal behaviors through interactions with environments. The capacity to continuously adapt and learn from new tasks without starting from scratch is a defining characteristic of human learning Kudithipudi et al. (2022) and a desirable trait for RL agents deployed in the dynamic and unpredictable real world. To achieve this, the research of continual reinforcement learning (CRL, a.k.a. lifelong reinforcement learning) has emerged Rolnick et al. (2019); Wolczyk et al. (2022); Kessler et al. (2022). It extends traditional RL by empowering agents with the ability to learn from a sequence of tasks, preserving knowledge from previous tasks, and using this knowledge to enhance learning efficiency and performance on future tasks. CRL aims to emulate the human capacity for lifelong learning and represents an ambitious field of research that tackles the challenges of long-term, real-world applications by addressing problems of increasing scale, diversity, and non-stationarity Khetarpal et al. (2022).
Despite progress in CRL, the field continues to face a critical challenge: insufficient transfer of knowledge across diverse tasks. Sufficient knowledge transfer is essential for a robust CRL system, yet many existing methods struggle to achieve this, leading to suboptimal performance and a failure to fully exploit the benefits of continual learning (CL) Wang et al. (2023c). A primary reason for this shortfall is the neglect of the sophisticated hierarchical structure of human cognitive control, which is primarily located in the prefrontal cortex and is complemented by the cerebellum for precise action control, enabling high-level cognition for envisioning the future and planning D’Mello et al. (2020); Friedman and Robbins (2022). This structure enhances humans’ ability to tackle complex tasks, and crucially, transfer skills and high-level knowledge across significantly diverse tasks. In contrast, as shown on the left side of Figure 1, existing CRL methods often focus only on transferring low-level knowledge, such as sharing the policy network. This limits them to sequences of highly similar tasks with slight variations only in goals or parameters Kessler et al. (2022); Gaya et al. (2023), and results in ineffective transfer across diverse tasks.
To take advantage of abstract similarities between diverse tasks, it is essential to represent high-level knowledge effectively and then employ robust models capable of extracting and transferring this knowledge. Human language, with its inherent abstraction, is a natural fit for representing high-level knowledge. Furthermore, recent breakthroughs in large-scale language research have demonstrated the impressive reasoning and in-context learning abilities of the large language model (LLM) Zhao et al. (2023); Xi et al. (2023); Wang et al. (2023b). Based on these insights, we propose a new framework named Hi-Core (Hierarchical knowledge transfer for Continual reinforcement learning), which integrates the powerful capabilities of an LLM into the CRL paradigm to enhance high-level knowledge transfer and eventually improve learning performance.
The right side of Figure 1 illustrates the idea of Hi-Core. Hi-Core is structured in two layers: the high-level policy formulation and the low-level policy learning. At the high level, Hi-Core utilizes the powerful reasoning ability of a LLM to formulate a high-level policy, which is represented as a temporally extended sequence of goals. Each goal consists of a descriptive component and an intrinsic reward function that outlines the intermediate states the agent needs to reach. The high-level policy is evaluated based on feedback from the low-level learning process, allowing the LLM to refine its policy. Following this process, the high-level policy along with the corresponding information is stored in a policy library. Upon encountering a new task, Hi-Core can retrieve this policy from the library as a context of the LLM for enhancing high-level knowledge transfer. At the low level, Hi-Core employs goal-oriented RL to learn a policy that is guided by the formulated high-level policy. The verified policy can be stored in the policy library to reduce the risk of catastrophic forgetting when encountering old tasks. Experiments are conducted in simple yet challenging MiniGrid environments. The results demonstrate that Hi-Core significantly outperforms popular baselines in terms of overall performance and transfer performance. The main contributions of the work are as follows:
We proposed Hi-Core, a novel framework that leverages hierarchical knowledge transfer for CRL, which aligns with human cognitive control, enhancing the effectiveness and generalization of agents.
To the best of our knowledge, Hi-Core is the first work to integrate the powerful reasoning ability of LLM into the CRL paradigm, facilitating high-level policy formulation and knowledge transfer across diverse tasks.
Extensive experiments in MiniGrid provide empirical evidence of Hi-Core’s effectiveness. Specifically, the results indicate that Hi-Core can be applied to heterogeneous tasks, achieving comparable transfer performance.
2 Background
2.1 Preliminaries
The continual reinforcement learning process can be formulated as a series of related Markov decision processes (MDPs) . Each MDP represents a different task or problem instance that an agent needs to solve over its lifetime. Here, and denote the state and action space of task , respectively, while is the transition probability function, and is the reward function. At each time step, the learning agent perceives the current state and selects an action according to its policy with parameters . The agent then transitions to the next state and receives a reward . The target of an agent on the task is to maximize the expected return , where is the discount factor of the reward, and is the horizon.
2.2 Related works
2.2.1 Continual Reinforcement Learning
CRL focuses on training RL agents to learn multiple tasks sequentially without prior knowledge, garnering significant interest due to its relevance to real-world AI applications Khetarpal et al. (2022). A central issue in CRL is catastrophic forgetting, which has led to various strategies for knowledge retention. PackNet and related pruning methods Mallya and Lazebnik (2018); Schwarz et al. (2021) preserve model parameters but often require knowledge of task count. Experience replay techniques such as CLEAR Rolnick et al. (2019) use buffers to retain past experiences, but face memory scalability challenges. In addition, some methods prevent forgetting by maintaining multiple policies or a subspace of policies Schöpf et al. (2022); Gaya et al. (2023). Furthermore, task-agnostic CRL research indicates that rapid adaptation can also help prevent forgetting Caccia et al. (2023).
Another issue in CRL is transfer learning, which is crucial for efficient policy adaptation. Naive approaches, like fine-tuning, that train a single model on each new task provide good scalability and transferability but suffer from catastrophic forgetting. Regularization-based methods, such as EWC Kirkpatrick et al. (2017); Wang et al. (2023c), have been proposed to prevent this side effect, but often reduce plasticity. Some architectural innovations have been proposed to balance the trade-off between plasticity and stability Rusu et al. (2016); Berseth et al. (2022). Furthermore, methods like OWL Kessler et al. (2022) and MAXQINIT Abel et al. (2018) leverage policy factorization and value function transfer, respectively, for improved learning across tasks.
Most existing methods perform well when applied to sequences of tasks that exhibit high environmental similarity, such as tasks where only specific parameters within the environment are altered or the objectives within the same environment are different. However, their effectiveness is greatly diminished when dealing with a sequence of diverse tasks. Our proposed framework aims to overcome this limitation by leveraging hierarchical knowledge transfer.
2.2.2 Reinforcement Learning with the LLM
Recent advancements have combined the large language model with reinforcement learning to address reinforcement learning challenges such as sample efficiency and generalization. These integrations utilize LLM’s vast knowledge and reasoning ability to improve agents’ performance. For instance, studies have developed methods for efficient agent-LLM interactions Hu et al. (2023) and unified agent foundations with language models for better experience reuse Palo et al. (2023).
Frameworks that merge LLMs with RL have also emerged, such as LAMP, which pretrains RL agents with Vision-Language Models Adeniji et al. (2023), and HiP, which uses foundation models for long-horizon tasks Ajay et al. (2023). In addition, the RAFA framework offers a principled approach with guarantees to optimize reasoning and actions Liu et al. (2023). Research into LLM-guided skill learning and reward shaping is also gaining traction. The BOSS method Zhang et al. (2023) and the work on LLM-based reward structures for robotics Yu et al. (2023) exemplify this trend, with studies like Kwon et al. (2023) focusing on LLM-driven reward design.
Our proposed framework extends these studies to continual learning in RL by leveraging LLM’s powerful reason ability for high-level policy formulation, thus enhancing high-level knowledge transfer.
3 Method
In this section, we clarify the research problem and describe the framework.
Problem Formalization. We formalize the CRL problem in the goal-oriented RL framework Chen and Luo (2022); Colas et al. (2022). The agent perceives not only the current state of the environment but is also provided with a specific goal , where represents the space of possible goals. For task and its corresponding goal sequence , the agent’s objective is to learn a policy that effectively maps each state to an appropriate action . The aim is to maximize the expected cumulative reward over time, guided by the sum of the extrinsic reward and the intrinsic rewards Dilokthanakul et al. (2019), where represents the next state in the task . The former is determined by the environment and is usually sparse. The latter is typically defined so that the agent receives a reward if it achieves the given goal.
Overview. Hi-Core is illustrated in Figure 2. For simplicity, we have omitted the task-indexed superscript in the description. The framework orchestrates the interaction between the CRL agent and the environment simulator for each task in the task sequence, aligning with standard RL protocols. For each task, the corresponding environment simulator provides the current state to the agent at each timestep, which then selects an action to perform, resulting in a reward and the next state. Internally, the CRL agent is structured in two layers: the high-level policy formulation and the low-level policy learning. Additionally, a policy library is maintained for the storage and retrieval of policies. The formulation of the high-level policy is initiated when the agent interacts first with the environment. If the policy library is not empty, related experiences (including high-level policies, corresponding environment descriptions and feedback) will be retrieved through the description of the environment to assist in the formulation. Once the high-level policy is established, it guides the learning of the low-level policy. Provided that there is sufficient storage, the most similar low-level policy can be retrieved from the policy library to serve as an advantageous initial point. Through feedback from the RL learning process, the LLM gradually refines its high-level policy to improve performance. Upon completion of a task, the successful policies are stored in the policy library.
3.1 High-level Policy Formulation
To integrate LLM into the goal-oriented RL framework, we conceptualize LLM as a function , where is the set of all prompts, is the set of textual descriptions of the environment, and is the set of textual descriptions of the sequence of goals. The description of the environment can be automatically obtained through a hard-coded reporter Dasgupta et al. (2022). The prompt is a concatenation of two parts: and , where is the system prompt and is the experience retrieved from the policy library. Description of each high-level policy, that is, textual output is mapped to the corresponding goal sequence through a defined parsing function. In order to improve the stability of LLM’s output, we structure the goal into a tuple: , where is the textual description of the goal state, is a verification function , is a intrinsic reward value after achieving the goal. Optional validation functions are provided in .
Intrinsic reward can not only guide the agent to learn to achieve the corresponding goal but also serve as a metric for evaluating whether the agent has achieved the goal. This metric is continuously monitored during the interactions of the CRL agent with the environment. When this average exceeds a predefined threshold, it is taken as an indication that the agent has successfully achieved the goals. The iterative process of high-level policy formulation, validation, and improvement draws inspiration from the concept of Reflexion Shinn et al. (2023). In instances where the agent fails to meet the threshold, indicating that the goal has not been met, the LLM is prompted to adjust the high-level policy. This adjustment is informed by feedback from the agent’s performance. The LLM first analyzes the feedback to pinpoint the reasons behind the shortfall in the previous policy. Then it leverages the powerful reasoning ability to hypothesize potential solutions. With this insight, the LLM formulates a new high-level policy aimed at overcoming the identified challenges.
3.2 Low-level Policy Learning
The low-level policy learning is guided by the formulated high-level policy. The RL model learns the low-level policy using raw state input and total intrinsic reward based on the goal sequence . The agent’s intrinsic reward function is defined as:
This reward structure provides a more granular view into the effectiveness of goal-related actions, encouraging the agent to learn a policy that is aligned with the goal more efficiently. Furthermore, the state-dependent design of this function improves the generalizability of high-level policies. Then, the internal rewards for each goal and the external reward from the environment will be textual and will be concatenated as the feedback. The total intrinsic reward is . Therefore, the aim of RL agent with goal sequence can be expressed as
3.3 Policy library
Recent studies Wang et al. (2023a); Zhang et al. (2023) have highlighted the potential of LLMs’ reasoning and planning ability in facilitating CL for agents, which involves continuous acquisition and update of skills Wang et al. (2023c). One method to address the core challenges in CL is to design a knowledge base Li et al. (2023) or a skill library Wang et al. (2023a). By using validated successful experiences, agents can not only reduce catastrophic forgetting but also rapidly enhance their capabilities.
Based on the above findings, Hi-Core maintains a policy library to store the policies that solve tasks successfully. The description of each high-level policy and the corresponding feedback is indexed by the embedding of the environment description of its corresponding task, which can be retrieved in similar tasks in the future. The policy library can be continuously extended throughout the life of the CRL agent. Given a new environment description , Hi-Core use embedding to retrieve the first similar environment descriptions and concatenate them with goal descriptions and corresponding feedback as experience . In addition, for each environment description , Hi-Core also stores the parameter of the low-level policy that has been learned. This can accelerate learning on similar tasks and further reduce catastrophic forgetting if storage space is sufficient. By storing both high-level and low-level learning policies in the policy library, the agent’s capabilities for diverse tasks can be enhanced over time.
4 Experiments
In this section, we evaluate our framework in several continual reinforcement learning tasks.
4.1 Environment
For our experiments, we utilized a suite of MiniGrid environments Chevalier-Boisvert et al. (2023) to evaluate the efficacy of Hi-Core in addressing CRL tasks. These environments feature image-based state observations, a discrete set of possible actions, and various objects characterized by their color and type. We intentionally simplified the state and action spaces to streamline the policy learning process, enabling more rapid experimental iterations. This approach allowed us to focus on the core objective of this study: leveraging hierarchical knowledge transfer for CRL. We crafted a sequence of four distinct tasks within the MiniGrid framework. Visual representations of each task environment are provided in Figure 3. This sequence is arranged in ascending order of difficulty and aligns with the human learning process, facilitating better tracking of knowledge transfer throughout the learning process.
4.2 Experiment Setup
Implementation. We use OpenAI’s GPT as the LLM, which can be accessed through their API. To improve the language model’s reasoning ability, we introduce the Chain of Thought (COT) Wei et al. (2022) in the prompt. This approach also aids the LLM in addressing complex tasks. Moreover, we add a few-shot example at the end of the prompt to guide the output towards the desired format. Note that this example contains only the foundational knowledge of all tasks, requiring the LLM to engage in additional reasoning to comprehend the specifics of the current task. For the low-level policy learning in Hi-Core, we chose the Proximal Policy Optimization (PPO) algorithm as the underlying RL algorithm due to its excellent and stable performance Schulman et al. (2017); Raffin et al. (2021). For the policy library, we build a vector database based on FAISS, which is a library for efficient similarity search and clustering of dense vectors Johnson et al. (2019). Due to the small total number of tasks, the number of experiences to be retrieved can be set as the total number of tasks . In addition, the policy library in our experiments does not store low-level policies for fair comparison with limited storage space.
Baselines. We consider the following baselines, for comparison with our approach in experiments:
Single-Task (SG): This baseline represents a traditional RL setup where a distinct agent is trained exclusively on each task. This serves as a foundational comparison point to underscore the advantages of CL, as it lacks any mechanism for knowledge retention or transfer.
Fine-Tuning (FT): Building upon the standard RL algorithm, this baseline differs from SG by using a single agent that is sequentially fine-tuned across different tasks. This baseline provides a basic measure of an agent’s capacity to maintain knowledge of earlier tasks while encountering new tasks.
Note that our method is not directly comparable with certain methods, such as CLEAR Rolnick et al. (2019) and ClonEx-SAC Wolczyk et al. (2022), as they require access to the data from previous tasks for training. Our framework can continually learn without the need to revisit past task data, which is a common constraint in real-world applications.
Metrics. To evaluate the effectiveness of Hi-Core, we use the normalized average return to measure the performance of the trained agents. Following standard practice in supervised continual learning literature Díaz-Rodríguez et al. (2018); Wolczyk et al. (2021), we define a suite of metrics based on the agent’s performance throughout different phases of its training process. These metrics include average performance , forward transfer and forgetting .
4.3 Competitive Results
To evaluate our framework, we compare Hi-Core with other baselines across four challenging tasks that have the same state and action space (Figure 3). The underlying RL algorithm of all baselines is PPO. Each experiment is trained in 5M steps and replicated with five random seeds of environments to ensure reliability.
Overall Performance. As illustrated in Figure 4, the baselines perform well on the first Task. However, their performance gets worse progressively as the difficulty of the tasks increases. A common trend of performance degradation in the transition between tasks is evident across all methods. Specifically, the SG baseline, while effective on the first task, is unable to maintain its performance due to its inherent limitation. Other baselines, such as FT, FT-L2, and PackNet, also start strong on the first task but encounter difficulties in the subsequent tasks. This phenomenon suggests that a singular shared policy may not be robust enough to handle a sequence of highly varied tasks. The balance between policy stability and plasticity is critical, with higher stability often leading to reduced adaptability, as observed in the performance degradation of these baselines. In contrast, Hi-Core consistently surpasses the other baselines, showcasing remarkable performance across all tasks. Although it is not immune to performance degradation when transitioning between tasks with significant differences, it demonstrates a rapid recovery in performance. This can be attributed to its hierarchical structure, which leverages the powerful capabilities of LLM and promotes the transfer of high-level policy knowledge, thereby enhancing the agent’s ability to adapt to new and diverse tasks.
| Metric | SG | FT | FT-L2 | PackNet | Hi-Core |
| 0.65 | |||||
| 0.17 | |||||
| 0.05 |
Continual Learning Performance. Table 1 shows the evaluation results in terms of CL metrics. The average performance metric clearly demonstrates Hi-Core’s superiority in task sequence with diverse tasks, significantly outperforming other baseline methods. The forward transfer metric is particularly important, as it measures the capacity of an agent to utilize knowledge from previous tasks. All baselines, with the exception of Hi-Core, present forward transfer metrics of less than or equal to zero. This suggests that traditional approaches struggle to transfer previously acquired knowledge to new and diverse tasks effectively. Baselines that rely on regularization strategies, such as FT-L2, or parameter isolation techniques, like PackNet, may even hinder the learning of new tasks.
In contrast, Hi-Core exhibits robust transfer performance, underscoring the benefits of transferring high-level policies. When examining the forgetting metrics, FT, FT-L2, and PackNet show relatively low scores. This can be attributed, in part, to their emphasis on model stability. However, a more critical factor is their inherently lower performance on subsequent tasks, such as Task 2, Task 3, and Task 4. Although Hi-Core does not achieve the lower score in the forgetting metric, its exceptional forward transfer capability and strong average performance accentuate its proficiency in handling diverse task sequences. Note that Hi-Core does not resort to fine-tuning strategy in this experiment, although it could be employed in situations where minimizing forgetting is a priority. The above findings highlight Hi-Core’s potential to markedly enhance the adaptability and longevity of reinforcement learning agents, positioning it as a highly viable solution for dynamic and diverse tasks.
| SG | |||
| Hi-Core | 0.65 | 0.17 | |
| Hi-Core-FT | 0.09 | ||
| Hi-Core-SG | |||
| Hi-Core-Once |
4.4 Heterogeneous Tasks Experiments
Most existing works have focused on the transfer ability of CRL agents in environments that have the same state space. In contrast, our study explores the application of CRL agents to heterogeneous tasks, where environments have different state spaces or action spaces. For the simplification of the problem, we only concentrate on tasks with differing state spaces.
As depicted in Figure 5a, heterogeneous tasks present a unique challenge. While it is possible to accommodate differences in the size of the state space during knowledge transfer by employing a masking technique, this approach necessitates prior knowledge of the maximum state space size, which is impractical for RL agents expected to learn continually in dynamic environments. On the contrary, the high-level policy transfer of Hi-Core can be applied to heterogeneous tasks without the need for masking.
In our experiments, we evaluate the effectiveness of Hi-Core across two heterogeneous tasks. The first task has a relatively small state space, as shown in the left below of Figure 5a. In contrast, the second task has a larger state space, such as the right below of the same figure. Figure 5b reports the results of these experiments. The first task is Task 1, while the second task is Task L1, Task L2, Task L3, or Task L4, each representing a larger state space variant of Task 1 through Task 4. Direct learning agents who are exposed to tasks with large state spaces from the outset may struggle to learn effective policies within the set number of steps, resulting in poor performance. This phenomenon is more obvious on more difficult tasks, such as Task L3 and Task L4. However, agents that are learned on Task 1 can transfer high-level knowledge to subsequent tasks. Despite the increasing difficulty of these tasks, the agents still demonstrate comparable performance. This evidence suggests that Hi-Core retains transfer ability even when applied to heterogeneous tasks.
4.5 Ablation Study
In this part, we conduct an ablation study to investigate what affects Hi-Core’s performance in CRL. We consider several variants of Hi-Core to understand the effects:
Hi-Core-FT: Hi-Core fine-tuning the policy from the previous task when encountering a new task.
Hi-Core-SG: Hi-Core without policy library. This variant treats each task as an isolated learning problem, akin to training single agents for each task.
Hi-Core-Once: Hi-Core without feedback mechanisms. In this variant, the high-level policy formulation of LLM is performed only once for each task.
The results of our study are presented in Table 2 and Figure 6. For better comparison, we also include the performance of the SG baseline in the Table. Our findings are as follows: Firstly, Hi-Core-SG shows superior performance compared to the SG baseline, which confirms that the LLM’s powerful capabilities can help RL. An RL agent can take advantage of reasoning ability in the LLM to significantly reduce the difficulties in solving these tasks. Secondly, while Hi-Core-SG outperforms the baseline, it falls short of the average performance and forward transfer metrics achieved by the Hi-Core. As shown in Figure 6b, this gap emphasizes the value of the policy library in facilitating knowledge transfer, ultimately boosting the agent’s overall performance. Furthermore, Hi-Core-Once exhibits slightly lower performance on Task 2, Task 3 and Task 4 than Hi-Core. This phenomenon suggests that feedback can improve the performance of LLM in dealing with these challenging tasks. Lastly, the Hi-Core-FT variant underperforms on all tasks except Task 1. This observation aligns with the previous finding in competitive results that methods designed to enhance policy stability may decrease the agent’s performance on subsequent tasks when those tasks differ significantly from one another.
Policy Visualization. Figure 7 further illustrates the difference between Hi-Core and Hi-Core-SG on Task 2 and Task 4. As shown in Figure 7a, the agent of Hi-Core demonstrates the capability to swiftly learn to pick up the blue key and unlock the corresponding door. In contrast, the agent of Hi-Core-SG struggles to progress beyond the initial room (Figure 7b). This performance gap can be attributed to the previous learning of the Hi-Core agent on Task 1, which shares common elements with Task 2, such as the blue key and door. By drawing on previous experience, Hi-Core formulate relevant initial goals, accelerating learning on Task 2. Similarly, when comparing Figure 7c and Figure 7d for Task 4, we observe a consistent pattern. Although Task 4 presents a higher level of difficulty, the agent with pre-acquired high-level knowledge exhibits more effective behaviors than the agent learning from scratch. These findings demonstrate the effectiveness of Hi-Core in facilitating the transfer of knowledge across diverse tasks.
5 Conclusion
In this work, we have presented Hi-Core, a novel framework that aligns with the hierarchical structure of human cognitive control, enhancing the high-level knowledge transfer of CRL. Hi-Core leverages a LLM for high-level policy formulation, employs goal-oriented RL for low-level policy learning, and constructs the policy library to store and retrieve policies. It addresses the challenge of learning across a sequence of diverse tasks by transferring high-level knowledge, which has been demonstrated in our experiments. Further results indicate that Hi-Core can be applied to heterogeneous tasks and achieves comparable transfer performance.
We believe that this work constitutes the first step toward the use of hierarchical knowledge transfer to enhance the effectiveness and generalization of CRL agents. There are still some limitations of Hi-Core, such as an unsatisfactory solution to catastrophic forgetting and room for improvement in the transferability of low-level policies. These directions provide opportunities for future work, with the potential to refine and extend Hi-Core.